# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v2

@@ III.3.3 Benchmarks in NISQ Era
We have introduced the main proposals of metrics and benchmarks in the literature. Table 1 analyze metrics, evaluating which quality attributes they fulfill according to the framework introduced in Sec. II. We emphasize that we evaluate the quality attributes fulfilled by metrics in the NISQ era; Therefore, these evaluations are not valid for the post-NISQ era. We hope that as quantum processors evolve, benchmarks will adapt to each era.
@@
(Average) Gate Fidelity $\times$ -- $\times$ Quality
(Average) Decoherence Times $\times$ -- $\times$ Quality
@@
Cross Entropy Benchmarking (XEB) $\times$ -- Quality
Hellinger distance $\times$ -- Quality
Heavy Output Generation $\times$ -- Quality
$\ell_{1}$-Norm Distance $\times$ -- Quality
Collision volume $\times$ -- Quality
Table 1: Summary of quantum computing metrics in NISQ era. In this table, we show some of the main metrics that have been proposed to measure properties related to the performance of a quantum computer, as presented in subsection III.3.2. We assess whether these metrics satisfy the quality attribute proposed in the literature for an ideal performance metric, as discussed in subsection II.3, although some of them are borderline. The check mark indicates that the metric satisfies the attributes; otherwise, we use the cross mark. The dash mark indicates that it is complicated to define if the metric satisfies the criterion because we cannot characterize it formally. Performance definitions: Scalability refers to the size of the processor; Quality evaluates how close the quantum processor is to a fault-tolerant quantum processor; Speed concerns the time required to process a certain task.
@@ IV.1 Guidelines for the benchmark practitioner
Guideline 1. Benchmarks must uphold the key quality attributes: relevance, reproducibility, fairness, verifiability, and usability. In Subsec. II.1, we introduced a set of quality attributes initially suggested as guidelines for classical benchmark designers. We aim to extend the application of these attributes to the field of quantum computing devices. Relevance aids in defining the specific area within the total behavior space targeted by our benchmark proposal and clarifies how accurately it captures this domain. Fairness is key for ensuring impartial comparisons across diverse gate-sets and processor architectures. Reproducibility and usability guarantee that results obtained by a particular device running the benchmark can be verified by other vendors, fostering trust among stakeholders. Finally, verifiability is instrumental in confirming the correctness of the results. A benchmark should be self-verifiable, to enable other parties to repeat the executions and obtain similar results.
@@
Guideline 2. Benchmarks must offer peak and base performance results. The classical baseline debate with arguments against and for an extended use of optimization flags can be extended to quantum devices. We encourage benchmarks practitioners to follow a similar approach as the final compromise established for SPEC in January 1994: ”Every SPEC CPU benchmark measurement has to measure performance with a restricted baseline set of flags, and optionally with a more extended peak set of flags”SPECHPGBenchmarksHistory . Having both metrics, peak and base, allows different vendors to share a common baseline for comparison that represents also the average non expert consumer, while enabling also the publication of the peak performance with optimised results leveraging expert knowledge over the architectures.
Guideline 3. Different technological eras necessitate different types of benchmarks. In the NISQ era, where there is no dominant architecture for quantum devices, specification-based benchmarks are particularly relevant. Given the substantial impact of noise, focusing on synthetic, microbenchmarks or kernel benchmarks could offer more practical and useful insights. As technology involves, Partially Corrected and Fault Tolerant quantum computers may transition into hybrid or kit-based benchmarks. Ultimately, for Fault Tolerant and Fully Functional Quantum Computers eras, application-based benchmarks would emerge as a reasonable and effective choice. As quantum computers progress, benchmarks will evolve to reflect this.
@@
Guideline 4. Benchmarks metrics must uphold the key quality attributes: practical, repeatable, reliable, and consistent. They should also preferably be independent and linear. As discussed in Sec. II, for a metric to be useful in build a good benchmark it must present attributes that correlate later with quality benchmark attributes. Additionally, while linearity and independence are desirable metric attributes, they are not strictly required for effectively capturing the performance of a quantum device.
@@
Guideline 5. Do not rely solely on a single benchmark to measure performance. While benchmarks serve as valuable tools for modeling specific aspects of the behavior of a quantum computer, relying solely on individual benchmarks may result in a lack of depth. The risk of spreading benchmarks too thin clearly arises when attempting to cover the entirety of the behavior of a device space with a single metric. Different behaviors often require distinct testing and experimentation setups, and consequently, an approach based on benchmark suites is recommended. Benchmark suites provide a more comprehensive assessment, enabling practitioners to really grasp how the computers perform in different situations.
@@ IV.2 Benchmarking structure
After completion of the benchmark execution and verification, we proceed to report the results. We must elaborate a reports compiling all results from the execution of the test suites for the peak and base configuration. Additionally, it should include the static and non-static metrics, the verification protocol, the configuration used for peak and base executions and execution process, so that others users can reproduce and verify your results. This report should be clear and contain as much information as possible so that users know how the computers are performing and can identify areas where the computers are underperforming.
@@
  • Definition of good metrics. In this study, we have introduced in Subsec. II.3 the quality attributes that good metrics should ideally meet. Additionally, we have examined various proposals put forth by researchers in the field in Subsec. III.3. However, we contend that further work is necessary to precisely define the relevance of the metrics. In other words, we need to determine which aspects of the devices behavior space are being captured by these measurements. Establishing a correlation between metrics and their actual relevance to performance is crucial for obtaining diverse perspectives of the device, and a thorough assessment. Throughout the NISQ era, prioritizing metrics centered on scalability, accuracy, and quality could drive the advancement of more robust and resilient devices, as speed currently does not hinder experimentation. However, speed may become a necessary performance indicator as we transition towards the Fault Tolerant era or implement additional partial quantum error correction methods.
@@ IV.3 Standard Performance Evaluation for Quantum Computers (SPEQC)
Leveraging the lessons learned and inspired by the creation of Standard Performance Evaluation Corporation (SPEC) in classical processor benchmarking, we propose the creation of an organization dedicated to Standard Performance Evaluation for Quantum Computers (SPEQC). Drawing inspiration from the founding principle of SPEC, ”let us no longer use small, synthetic programs as benchmarks, let us use real application programs” SPECHPGBenchmarksHistory , should aim to define metrics similar to the SPECmark. The proposed metrics should avoid replicating what was done with MIPS (Millions of Instructions Per Second) in the 1980s. At that time, MIPS was recognized as an inaccurate measure of actual computer performance, leading to the development of more appropriate metrics. Similarly, SPEQC’s metrics should aim to replace current efforts to establish a standard metric to measure the performance of quantum computers.
@@
The development of benchmarks is a complex task and requires making significant decisions that must satisfy several design objectives simultaneously. In other words, when designing benchmarks, it is necessary to balance different objectives that are important for the design and functionality of the benchmark. These objectives may include easy to use, relevance, and fairness, among others. One of the most relevant aspects to highlight is the need to ensure that the benchmarks developed are fair. To achieve this, SPEQC should actively involve and collaborate with a diverse range of stakeholders. This would involve participation from representatives various sectors, including the scientific community, industry, public administration, and hardware and software vendors. A benchmark developed with the participation of all stakeholders is considered fairer than one designed exclusively by a single company. Although involving all stakeholders may not be the most efficient method for developing a benchmark, consensus among multiple stakeholders will make the benchmark fair. Consequently, SPEQC should adopt an inclusive membership policy, whereby any company or organization with an interest in advancing the field of quantum computing is encouraged and welcomed to join the initiative. This inclusive approach would ensure that the development and implementation of benchmarks are reflective of the broader interests and needs of the quantum computing community, thereby maximizing their impact and relevance in the industry. By promoting an open and collaborative environment for debate, the exchange of knowledge, perspectives, and best practices is facilitated, driving innovation and progress in the field of quantum computing.

### v3

@@ III.3.3 Benchmarks in NISQ Era
We have introduced the main proposals of metrics and benchmarks in the literature. Table 1 analyzes metrics, evaluating which quality attributes they fulfill according to the framework introduced in Sec. II. We emphasize that we evaluate the quality attributes fulfilled by metrics in the NISQ era; Therefore, these evaluations are not valid for the post-NISQ era. We hope that as quantum processors evolve, benchmarks will adapt to each era.
@@
Gate Fidelity $\times$ - $\times$ Quality
(Average) Decoherence Times $\times$ - $\times$ Quality
@@
Cross Entropy Benchmarking (XEB) $\times$ - Quality
Hellinger distance $\times$ - Quality
Heavy Output Generation $\times$ - Quality
$\ell_{1}$-Norm Distance $\times$ - Quality
Collision volume $\times$ - Quality
Table 1: Summary of quantum computing metrics in NISQ era. In this table, we show some of the main metrics that have been proposed to measure properties related to the performance of a quantum computer, as presented in subsection III.3.2. We assess whether these metrics satisfy the quality attribute proposed in the literature for an ideal performance metric, as discussed in subsection II.3, although some of them are borderline. The check mark indicates that the metric satisfies the attributes; otherwise, we use the cross mark. The dash mark indicates that it is complicated to define if the metric satisfies the criterion because we cannot characterize it formally. Performance definitions: Scalability refers to the size of the processor; Quality evaluates how accurately the quantum processor performs the tasks it was designed or programmed to execute; Speed concerns the time required to process a certain task.
@@ IV.1 Guidelines for the benchmark practitioner
Guideline 1. Different technological eras necessitate different types of benchmarks. In the NISQ era, where there is no dominant architecture for quantum devices, specification-based benchmarks are particularly relevant. Given the substantial impact of noise, focusing on synthetic, microbenchmarks or kernel benchmarks could offer more practical and useful insights. At this early stage of hardware development, it is crucial that benchmarks not only produce comparative scores but also provide detailed information about the performance of individual hardware components. This allows hardware developers to focus on improving their platforms rather than competing solely for benchmark scores. If a benchmark does not provide useful insight to help hardware developers identify which components hinder performance, it does not contribute meaningfully to the improvement of the device. In this context, benchmarks should be understood as tools for guiding technical progress, particularly during the early stages of quantum hardware. This does not mean that other types of benchmarks should not be excluded. In contrast, it is important to design application-oriented benchmarks for algorithms that are relevant to the current era, such as quantum machine learning (QML) algorithms or variational quantum algorithms (VQAs). However, it is less meaningful to prioritize the design of benchmarks that target applications that are not yet feasible with current quantum hardware, such as Shor or Grover algorithms, which are more relevant to future eras of quantum computing. However, recent work has proposed benchmarks based on these algorithms, indicating that this guideline, while seemingly obvious, is not universally followed. As technology involves, Partially Corrected and Fault-Tolerant quantum computers may transition into hybrid or kit-based benchmarks. In this Partially Quantum Error Corrected era we will also face the task of developing benchmarks for logical qubits and the quantum error correction methods applied to them, giving rise to a new class of benchmarks focused on fault-tolerant performance and error correction efficacy. Ultimately, for Fault Tolerant and Fully Functional Quantum Computers eras, application-based benchmarks would emerge as a reasonable and effective choice. As quantum computers progress, benchmarks will evolve to reflect this.
Guideline 2. Benchmarks must uphold the key quality attributes: relevance, reproducibility, fairness, verifiability, and usability. In Subsec. II.1, we introduced a set of quality attributes initially suggested as guidelines for classical benchmark designers. We aim to extend the application of these attributes to the field of quantum computing devices. Relevance aids in defining the specific area within the total behavior space targeted by our benchmark proposal and clarifies how accurately it captures this domain. Fairness is key for ensuring impartial comparisons across diverse gate-sets and processor architectures. Reproducibility and usability guarantee that results obtained by a particular device running the benchmark can be verified by other vendors, fostering trust among stakeholders. Finally, verifiability is instrumental in confirming the correctness of the results. A benchmark should be self-verifiable, to enable other parties to repeat the executions and obtain similar results. One of our key proposals and recommendations is that any newly defined benchmark should be accompanied by an analysis that clearly identifies which properties it satisfies and which it does not, along with a justification for the relevance or irrelevance of each property in the given context. This practice will help the community understand the appropriate contexts in which these benchmarks or metrics should be applied, as well as highlight key limitations that future developments should aim to address. As an illustrative example, we briefly analyze one of the benchmarks previously discussed in Subsec. III.3.3: the QED-C Benchmark. This benchmark, or benchmark suite, has been continuously updated to incorporate a variety of application-oriented benchmarks for quantum computers. The final output of the benchmark can be interpreted either as a vector comprising the results of multiple specified metrics, or as a collection of individual benchmarks, each with its own set of metrics. Regardless of the interpretation, the resulting analysis remains largely consistent. In this analysis, we treat it as a benchmark suite. Some components of the suite, such as those based on Shor or Grover algorithms, are not relevant in the current technological era, a point acknowledged in the original guidelines. Others, such as benchmarks based on the Variational Quantum Eigensolver (VQE), remain pertinent. In principle, the benchmarks are reproducible, provided that the corresponding metrics are reported alongside detailed information about all parameters and device calibration. They are also fair, as they do not introduce biases tied to specific hardware architectures or artificial constraints. The methodology is well-defined and established, contributing to the overall verifiability of the suite. However, some benchmarks, such as those involving Hamiltonian simulation, do not scale efficiently to larger qubit counts due to their reliance on classical simulation. This reliance renders them unverifiable for systems with many qubits, as classical resources become insufficient to reproduce or validate the results. Moreover, not all benchmarks in the suite are usable at scale; for example, in cases like Hamiltonian simulation, the required number of measurement shots increases significantly with the number of qubits, making them unsuitable for systems with many qubits.
@@
Guideline 3. Benchmarks must offer peak and base performance results. The classical baseline debate with arguments against and for an extended use of optimization flags can be extended to quantum devices. We encourage benchmarks practitioners to follow a similar approach as the final compromise established for SPEC in January 1994: ”Every SPEC CPU benchmark measurement has to measure performance with a restricted baseline set of flags, and optionally with a more extended peak set of flags”SPECHPGBenchmarksHistory . Having both metrics, peak and base, allows different vendors to share a common baseline for comparison that represents also the average non expert consumer, while enabling also the publication of the peak performance with optimised results leveraging expert knowledge over the architectures.
In classical computing, compilers are highly advanced and well-established, enabling standardized configuration through flags across different hardware architectures. In contrast, the quantum computing field lacks such standardization: each device typically relies on its own compiler, with no unified approach to transpilation or even to the overall process of circuit preparation. This fragmentation makes it challenging to define a consistent set of configuration flags. Nonetheless, it is possible to identify a set of relevant configuration parameters that could serve as flags in the current NISQ era. It is important to emphasize that each era of quantum computing requires its own definitions of baseline and peak performance. Some example configuration parameters include:
  • Classical subroutine simulation: Allow simulating parts of the circuit or specific subroutines on a classical computer when feasible. For instance, Clifford circuits can be efficiently simulated classically.
  • Circuit cancellation: Permit the elimination of subcircuits that have no effect on the final output, such as a unitary followed immediately by its inverse.
  • Circuit equivalence transformations: Allow transformations of the circuit into another that is equivalent under a well-defined criterion, e.g., preserving the final quantum state. For example, conjugating a Pauli gate with a Clifford circuit results in another Pauli gate.
  • Approximate circuit transformation: Permit transformations that yield a circuit producing a result $\epsilon$-close (under a defined metric) to the original target outcome. An example includes rearranging gates that do not strictly commute to obtain a similar final state.
  • Error mitigation: Specify whether error mitigation techniques are allowed, and if so, define which classes of techniques (e.g., zero-noise extrapolation, probabilistic error cancellation) are permissible.
  • Measurement strategies: Include flags for measurement techniques, which are especially relevant in the NISQ era due to the significant error contribution from the measurement process. Examples include changing the measurement basis to reduce the number of shots (e.g., grouping commuting Pauli observables) or using advanced techniques like classical shadows to efficiently estimate multiple observables.
  • Hardware optimization: Techniques that depend on the specific characteristics of the quantum device, such as qubit fidelity, coherence times, gate connectivity, native operations, among others. For example, insert the minimal number of SWAP gates when qubits are not directly connected, selecting the lowest-cost path according to the device connectivity.
@@
Guideline 4. Benchmarks metrics must uphold the key quality attributes: practical, repeatable, reliable, and consistent. They should also preferably be independent and linear. As discussed in Sec. II, for a metric to be useful in build a good benchmark it must present attributes that correlate later with quality benchmark attributes. Additionally, while linearity and independence are desirable metric attributes, they are not strictly required for effectively capturing the performance of a quantum device. Regarding benchmarks, we encourage that each proposal of a new metric includes an analysis of the properties it satisfies as well as those it does not.
@@
Guideline 5. Do not rely solely on a single benchmark to measure performance. While benchmarks serve as valuable tools for modeling specific aspects of the behavior of a quantum computer, relying solely on individual benchmarks may result in a lack of depth. The risk of spreading benchmarks too thin clearly arises when attempting to cover the entirety of the behavior of a device space with a single metric. For example, metrics such as Quantum Volume or Cross-Entropy Benchmarking (XEB) can provide a general overview of a quantum computer’s performance. However, to obtain a more comprehensive understanding, it is essential to incorporate additional, more specific metrics that assess individual characteristics of the system. Different behaviors often require distinct testing and experimentation setups, and consequently, an approach based on benchmark suites is recommended. Benchmark suites provide a more comprehensive assessment, enabling practitioners to really grasp how the computers perform in different situations.
@@ IV.2 Benchmarking structure
After completion of the benchmark execution and verification, we proceed to report the results. We must elaborate reports compiling all results from the execution of the test suites for the peak and base configuration. Additionally, it should include the static and non-static metrics, the verification protocol, the configuration used for peak and base executions and execution process, so that others users can reproduce and verify your results. This report should be clear and contain as much information as possible so that users know how the computers are performing and can identify areas where the computers are underperforming.
@@
  • Definition of good metrics. In this study, we have introduced in Subsec. II.3 the quality attributes that good metrics should ideally meet. Additionally, we have examined various proposals put forth by researchers in the field in Subsec. III.3. However, we contend that further work is necessary to precisely define the relevance of the metrics. In other words, we need to determine which aspects of the devices behavior space are being captured by these measurements. Establishing a correlation between metrics and their actual relevance to performance is crucial for obtaining diverse perspectives of the device, and a thorough assessment. Throughout the NISQ era, prioritizing metrics focused on scalability, accuracy, and quality can better guide the development of robust and practically useful quantum devices. While gate speed remains a relevant and informative metric—especially given the variability in gate durations across different technological platforms—we consider it complementary to those assessing overall system performance and reliability. In our view, scalability, accuracy, and quality represent more immediate challenges for current quantum computing architectures. Therefore, speed-related metrics are valuable, but should be regarded as secondary to those that capture the system’s computational potential and robustness. However, speed may become a necessary performance indicator as we transition towards the Fault Tolerant era or implement additional partial quantum error correction methods.
@@ IV.3 Standard Performance Evaluation for Quantum Computers (SPEQC)
While the development of quantum computing standardized benchmarks is a valuable and necessary effort, it is important to recognize that a standardization may introduce certain risks. One of these risks is that early standardization of benchmarking could unintentionally hinder the advancement of promising technologies that are still in earlier stages of development. This risk arises when benchmarking criteria reflect only the capabilities of more mature quantum hardware, thereby creating implicit performance expectations that emerging technologies cannot meet. As a result, research and investment may become concentrated on a narrow subset of quantum platforms, reducing diversity and potentially overlooking technologies with long-term advantages. To avoid this outcome, it is essential that benchmarking frameworks be flexible, inclusive, and responsive to the evolving technological landscape of quantum computing. For example, benchmarks could be designed in a modular way to provide flexibility, adopting layered structures, such as synthetic, microbenchmarks, kernel, hybrid, kit-based and application-level benchmarks, that can be tailored to the current stage of the technology under evaluation (NISQ, PQEC, FTQC, FFQC). To move in this direction, it would first be necessary to classify and characterize both existing and future benchmarks, and subsequently to accurately determine the technological stage of the quantum platform under evaluation. This would enable the application of appropriate criteria and fair evaluation based on the hardware level of maturity.
Building on these ideas and drawing inspiration from the establishment of the Standard Performance Evaluation Corporation (SPEC) in classical processor benchmarking, we propose the creation of an organization dedicated to Standard Performance Evaluation for Quantum Computers (SPEQC) as a flexible benchmarking framework that adapts to the different stages and technologies of quantum hardware development. Rather than enforcing a fixed set of metrics or algorithms, SPEQC should be designed to accommodate a diverse range of benchmark types. It is worth noting that SPEQC should not be limited to application or full-stack benchmarks. Instead, it should also incorporate low-level methodologies, including synthetic, microbenchmark, and kernel-based approaches, thereby providing a comprehensive framework for evaluating both individual hardware components and overall system performance across diverse quantum architectures.
@@
The development of benchmarks is a complex task and requires making significant decisions that must satisfy several design objectives simultaneously. In other words, when designing benchmarks, it is necessary to balance different objectives that are important for the design and functionality of the benchmark. These objectives may include ease to use, relevance, and fairness, among others. One of the most relevant aspects to highlight is the need to ensure that the benchmarks developed are fair. To achieve this, SPEQC should actively involve and collaborate with a diverse range of stakeholders. This would involve participation representatives from various sectors, including the scientific community, industry, public administration, and hardware and software vendors. A benchmark developed with the participation of all stakeholders is considered fairer than one designed exclusively by a single company. Although involving all stakeholders may not be the most efficient method for developing a benchmark, consensus among multiple stakeholders will make the benchmark fair. Consequently, SPEQC should adopt an inclusive membership policy, whereby any company or organization with an interest in advancing the field of quantum computing is encouraged and welcomed to join the initiative. This inclusive approach would ensure that the development and implementation of benchmarks are reflective of the broader interests and needs of the quantum computing community, thereby maximizing their impact and relevance in the industry. By promoting an open and collaborative environment for debate, the exchange of knowledge, perspectives, and best practices is facilitated, driving innovation and progress in the field of quantum computing.